OpenStudy (anonymous):

8 children have 9 candy. In how many ways we can divide candy for children? (one children has to have at least one candy) (candy are all different and children are all different)

OpenStudy (anonymous):

Assigning the first \(8\) pieces is just \(8!\). Then to assign the last piece there are \(8\) more choices. This gives us \(8!\times8\)

OpenStudy (anonymous):

Tasks done in sequence are multiplied.

OpenStudy (anonymous):

but what if I take, for example, 2 children and 3 candy and, if i write down every possibility i get 6, but using your example it should be 2! x 2

OpenStudy (anonymous):

Are you sure? Can I see your possibility list?

OpenStudy (anonymous):

Let the candies be r, b, g and a comma separating for each kid...

OpenStudy (anonymous):

Oh I see the error now...

OpenStudy (anonymous):

It matters which candy the 9th candy is.

OpenStudy (anonymous):

Okay, so how about 9C8 to determine ways to pick which candies will be the first 8. Then 8! to assign those candies. Lastly 9C1 to pick 9th candy and 8 ways to assign it.

OpenStudy (anonymous):

Then for 3 candy 2 kids. it is \(^3C_2 \times 2! \times ^3C_1\times 2\)

OpenStudy (anonymous):

Hmmm I wonder...

OpenStudy (anonymous):

maybe it should be something like if we have 2 children and 3 candy, then all possibilities are 2^3, with possibilities that one of the children won't get any candy at all, so we have to also take away these 2 possibilities. it works with smaller numbers, but i don't know what about the rest...

OpenStudy (anonymous):

Nope that doesn't seem to work either.

OpenStudy (anonymous):

Then you can to figure out how to count how many ways a kid can not get candy.

OpenStudy (anonymous):

Okay we could do it like this...

OpenStudy (anonymous):

Actually never mind I was thinking something wrong.

OpenStudy (anonymous):

8^9 is all possibilities with no candy for somone?

OpenStudy (anonymous):

All possibilities... even that a kid gets NO candy.

OpenStudy (anonymous):

If there are 4 kids... how many ways do you count?

OpenStudy (anonymous):

Here is my new idea. We chose one of the 9 candy which will be the bonus candy. We then give that to one of the 8 kids. Then after that we forget about the ninth candy and assign the remaining 8 candy. So its: \((9\times 8) \times 8!\)

OpenStudy (anonymous):

i've counted for 2 children and 3 candy and 2 children and 4 candy and i've written them all down

OpenStudy (anonymous):

can you tell me your numbers?

OpenStudy (anonymous):

2 children and 3 candy are 6 possibilities and 2,4 are 14 possibilities

OpenStudy (anonymous):

I think I figured out my error

OpenStudy (anonymous):

Okay so we start off by lining up the kids and have each one draw a random candy. This is \(^9P_8\)

OpenStudy (anonymous):

Have you done 3 children 4 candy?

OpenStudy (anonymous):

no, i think it is too difficult to count, i should be less than 81 i think

OpenStudy (anonymous):

\[ a-bc\\ b-ac\\ c-ab\\ ab-c\\ ac-b\\ bc-a \]

OpenStudy (anonymous):

\[ a-b-cd\\ a-c-bd\\ a-d-bc\\ b-a-cd\\ b-c-ad\\ b-d-ac\\ \]

OpenStudy (anonymous):

I shouldn't have been using permutations I think.

OpenStudy (anonymous):

Hmmm we gotta split it up into simple tasks.

OpenStudy (anonymous):

Okay here is one idea... 1. Choose 1 kid (8) 2. Have the kid choose 2 candy (9C2) 3. Have the remaining kids chose their candies ((8-1)!)

OpenStudy (anonymous):

\[ 2\times \binom{3}{2}\times (2-1)! = 2\times 3\times 1 = 6 \]

OpenStudy (anonymous):

For 3-4: \[ 3\times \binom{4}{2}\times (3-1)! = 3\times 6 \times 2= 36 \]

OpenStudy (anonymous):

and what if we take 2-4?

OpenStudy (anonymous):

\[ 8\times \binom{9}{2}\times (8-1)! = 8 \times 36 \times 5040= 1451520 \]

OpenStudy (anonymous):

It doesn't work for 2-4 because the problem with 2-4 is that you have many ways of distributing the candy.

OpenStudy (anonymous):

I mean you can do 2-2, 1-3, 3-1

OpenStudy (anonymous):

My method isn't generalized to that extent. I only works for distributing n among n-1

OpenStudy (anonymous):

ok, i understand now, thank you :)

OpenStudy (anonymous):

Ummm I wish we had list of 3-4 to check.

OpenStudy (anonymous):

I'm getting 36 possibilities so it wouldn't be impossible to list them out.

OpenStudy (anonymous):

a.b.cd a.c.bd a.d.bc b.a.cd b.d.ac b.c.ad c.a.bd c.b.ad c.d.ab d.a.bc d.b.ac d.c.ad That's 12, when the third guy gets the candy. By symmetry you count for the other three to get \(3\times 12=36\)